at a constant temperature 250 ml of argon at 760mm pressure and 600 ml of nitrogen at 500 mm pressure are put together in a one litre flask . calculate final pressure

Grade:11

1 Answers

Amardeep Kumar
6 years ago
V1= 250 mlV2= 600 mlP1= 760 mmP2 = ?P1/P2 = V1/V2P2 = p1 × V2/V1P2 = 760×600/250P2 = 9120 mmI just try